Rhodium has an FCC crystal structure, an atomic radius of 0.1345 nm and atomic weight of 102.9 g/mol. Calculate the density of rhodium.


1
Expert's answer
2022-03-08T12:24:05-0500

"\rho" = (nAv) / (VcNA)

n - atoms in unit cell,

Av - atomic weight,

Vc - unit cell volume,

NA - Avogadro's number

In a FCC lattice there are 8 atoms at eight corners and 6 at face centers. Now each corner contributes to eight cells so per unit cell contribution is 1/8×8=1 atom. Now similarly each face center contributes to two unit cells so contribution per unit cell by six face centers is equal to 1/2​×6=3 atoms. Hence total number of atoms per unit cell of FCC lattice is n = (1+3) = 4 atoms.

Av = 102.9 g/mol

For FCC unit cell volume Vc = 16R3√2 = (16)(0.1345 × 10-9 m)(√2) = 5.5 × 10-29 m3 = 5.5 × 10-23 cm3

NA = 6.022 × 1023

"\rho" = [(4 atoms/unit cell)(102.9 g/mol)] / [(5.5 × 10-23 cm3)(6.022 × 1023 atoms/mol)] = 12.43 g/cm3